You are taking out a $29556 loan. It will be amortized with fixed payments over 10 years. It is to be paid quarterly and the APR is 4%.

What is the interest payment on the loan in the second quarter?
Answer #1

Loan Amount = $29,556

Time Period = 10 years

Interest Rate = 4% quarterly

Loan has quarterly Payments,

So,

Period = 4*10 = 40

Interest Rate = 1% per quarter

Calculating Quarterly Payment,

Using TVM Calculation,

PMT = [PV = 29556, FV = 0, T = 40, I = 0.01]

PMT = $900.15

So,

Principal Payment in Quarter 1 = 900.15 - (0.01)(29556) = $604.59

Principal Due = $28,951.41

Interest Payment in Quarter 2 = 0.01(28951.41) = $289.51

So,

Interest payment in Quarter 2 = $289.51

Please "Like" if you find this useful, and comment if you have any doubts.